OpenStudy (anonymous):

Which proportion models this problem? On a soccer team, 5 out of 13 players surveyed said they had two or more siblings. The league has 520 players. Which is the best prediction of the number of players in the league that have two or more siblings?

OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!!!!!!!!!

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

We have x/520 and we have 5/13, so we make these equal to each other to make a proportion.

OpenStudy (anonymous):

so what would it be ? @iGreen

OpenStudy (igreen):

I'm asking you :) Make x/520 and 5/13 equal to each other.

OpenStudy (anonymous):

would x=200 @iGreen

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

But it only asks for the proportion: \(\dfrac{x}{520} = \dfrac{5}{13}\)

OpenStudy (anonymous):

so what does that mean? @iGreen

OpenStudy (igreen):

It means, this is your answer: \(\dfrac{x}{520} = \dfrac{5}{13}\)

OpenStudy (anonymous):

okay thank you can you help me with one more? @iGreen

OpenStudy (anonymous):

A bottling machine filled 3 of every 8 bottles with grape juice. The bottling machine filled 144 bottles altogether. How many of these bottles were filled with grape juice?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

One question per post, close this one and open a new one. @ineedurhelpnow